Show complete solution:

1.f(x)=2 sin (x-​π/​4)

2.f(x)=sin (x+π/​3)

3.f(x)=3 cos (x+π/​6)

4.f(x)=2 cos (x-2π)

5.f(x)=tan x



1
Expert's answer
2021-03-23T06:34:31-0400

"\begin{array}{l}
\text { 1) } f(x)=2 \sin \left(x-\frac{\pi}{4}\right)=0 \\
\sin \left(x-\frac{\pi}{4}\right)=0, \quad y=x-\frac{\pi}{4} \\
\sin y=0 \\
k \in Z \quad y=\pi k \quad x-\frac{\pi}{4}=\pi k \\
x=\frac{\pi}{4}+\pi k
\end{array}"


"\begin{array}{l}
\text { d) } 4) f(x)=2 \cos (x-2 \pi)=0 \\
\cos (x-2 \pi)=0 \\
\cos y=0 \quad y=x-2 \pi \\
y=\frac{\pi}{2}+\pi k, \quad x=\frac{\pi}{2}+2 \pi+\pi k \quad k \in z \\
x=\frac{5 \pi}{2}+\pi k \Rightarrow x=\frac{\pi}{2}+\pi k
\end{array}" "\begin{array}{l}
\text { 3) } \begin{aligned}
f(x)=3 \cos (x+\pi / 6)=0 \\
\cos (x+\pi / 6)=0
\end{aligned} \\
\text { Ces } \begin{array}{rl}
y=x+\pi / 6 \quad \cos y=0 \\
y=\frac{\pi}{2}+\pi k \quad x=\frac{\pi}{2}-\frac{\pi}{6}+\pi k & k \in Z \\
x=\frac{\pi}{3}+\pi k
\end{array}
\end{array}" "\text { 5) } \begin{aligned}
f(x) &=\tan x=0 \\
x &=\frac{\pi}{2}+\pi k \quad k \in Z
\end{aligned}" "\text { 2)} \begin{array}{rl}
f(x)=\sin (x+\pi / 3)=0 \\
\sin y=0, & y=x+\pi / 3 \\
y=10 \pi k & x+\pi / 3=\pi k \quad k \in Z \\
x & =-\pi / 3+\pi k
\end{array}"